# WhiSPA: Semantically and Psychologically Aligned Whisper with Self-Supervised Contrastive and Student-Teacher Learning

### v1

@@ Abstract
Current speech encoding pipelines often rely on separate processing pipelines between text and audio, not fully leveraging the inherent overlap between these modalities for understanding human communication. Language models excel at capturing semantic meaning from text that can complement the additional prosodic, emotional, and acoustic cues from speech. This work bridges the gap by proposing WhiSPA [^1] https://github.com/humanlab/WhiSPA (Whisper with Semantic-Psychological Alignment), a novel audio encoder trained with a contrastive student-teacher learning objective. Using over 500k speech segments from mental health audio interviews, we evaluate the utility of aligning Whisper’s audio embeddings with text representations from an SBERT encoder and text-based assessments of psychological dimensions: emotion and personality. Over self-supervised and downstream mental health tasks, WhiSPA surpasses state-of-the-art speech models, achieving an average error reduction of $73.4\%$ on the segment-level self-supervised objective and $83.8\%$ on 11 psychological downstream tasks. WhiSPA demonstrates that cross-modal alignment can increase the amount of text-semantic and psychological information captured in audio-only encoder models.
@@
Rajath Rao1, Adithya Ganesan1, Oscar Kjell1, Jonah Luby1, Akshay Raghavan1, Scott Feltman1 Whitney Ringwald4, Ryan L. Boyd3, Benjamin Luft1, Camilo Ruggero3, Neville Ryant2, Roman Kotov1, H. Andrew Schwartz1, 1Stony Brook University, 2University of Pennsylvania 3University of Texas at Dallas, 4University of Minnesota Correspondence: rajath.rao@stonybrook.edu, has@cs.stonybrook.edu
@@ 1 Introduction
Figure 1: The goal of WhiSPA is to align Whisper’s (speech-based) representations to better capture the semantic and psychological characteristics of communications that are currently best captured from text-based language models. Our approach enables speech-based models to enrich audio representations, which already capture prosodic cues, with stronger semantic and psychological dimensions.
Human communication is inherently multimodal, combining semantic, prosodic, and psychological cues to convey meaning. However, AI integration of modalities is often fragmented Lazaro et al. (2021); Gu et al. (2017), limiting models’ ability to encode both acoustic and semantic dimensions. Speech models excel at capturing prosodic features but lack the nuanced semantic understanding of the text that language models provide Soubki et al. (2024); Sriram et al. (2017).
@@
To bridge this gap, we introduce an audio encoding model, WhiSPA (Whisper with Semantic and Psychological Alignment), by attempting to address the semantic and psychological representation disparity between Whisper-based representations and text-based LLMs. We align a pre-trained speech transcription model, Whisper Radford et al. (2022), with latent dimensions from SBERT Reimers and Gurevych (2019) and language-based models of psychological attributes V Ganesan et al. (2022); Park et al. (2014), intended to carry deeper semantic and psychological dimensions. Such alignment not only reduces computational and memory inefficiencies (from not needing to run a second LLM encoder on the transcripts from an audio model), but also enables a unified understanding of cross-modal dependencies between speech and language models which are often trained on much larger corpora.
While text-based language models effectively capture semantics and some degree of human affective information, audio inherently contains additional acoustic information. For example, prosodic elements such as tone of voice and pause duration can alter the meaning or better convey affective (i.e. as pertaining to human emotion or tone) aspects of language — transforming a statement into a question, conveying sarcasm instead of sincerity, or shifting a serious remark into insincere humour. Without cross-modal integration, language models, while semantically robust, remain incomplete in representing the full spectrum of human expressions Zhang et al. (2023); Lian et al. (2023). Still, since text is derivable from speech, speech should intrinsically be mappable to the same rich semantic embeddings from language models, making an additional encoder redundant. By addressing this gap, our work streamlines the speech analysis pipeline and enhances its usability.
Contributions.
We hypothesize that aligning text and audio latent spaces can significantly improve audio-based representations. We test this hypothesis and make the following specific contributions: (1) We propose and train a novel audio encoder, WhiSPA (Whisper with Semantic and Psychological Alignment), which demonstrates significantly greater performance on both speech-to-text self-supervised tasks and downstream psychological tasks; (2) We systematically evaluate variants on the proposed alignment architecture, finding (a) aligning with text-based representations of both semantics and psychological features drastically improves audio representations, (b) aligning with emotional and personality factors result in SotA person-level psychological assessments, (c) a contrastive learning objective results in a superior point of convergence for encoding semantic and psychological dimensions of audio, and finally, (d) for downstream tasks, we found marginal benefit in adding SBERT representations to WhiSPA, suggesting that WhiSPA already captures nearly all the information provided by its text-based teacher model.
WhiSPA integrates acoustics with language-derived semantics and psychological features, enabling a more context-aware multimodal AI system to encode semantic and psychological dimensions from speech.
@@ 2 Background
Transformer architectures have revolutionized natural language processing (NLP) Vaswani et al. (2017), enabling significant advances in speech processing and cross-modal learning by setting new benchmarks for understanding language and multimodal data.
Speech Recognition.
Whisper, OpenAI’s state-of-the-art automatic speech recognition (ASR) model, was pre-trained on $680,000$ hours of multilingual, multitask audio data. Leveraging an encoder-decoder architecture, it performs speech transcription and translation Radford et al. (2022). Its autoregressive decoder predicts the next most probable token while attending to prior tokens and cross-attending to the audio representations from the encoder. Whisper’s intermediary layers encode high-level audio representations, capturing both phonetic and semantic content, enabling contextual understanding beyond sound patterns Radford et al. (2022).
The encoder processes Mel filterbank representations into embeddings encapsulating rich acoustic and linguistic features, such as phonemes, intonation, and broader context. The decoder refines this understanding, generating text even in ambiguous or noisy conditions. This capability mirrors human speech perception, where phonetic accuracy and contextual understanding are seamlessly integrated Radford et al. (2022).
Audio Encoders.
Audio-specific models like Wav2Vec have advanced the encoding of audio signals into high-dimensional representations. Wav2Vec Schneider et al. (2019), developed by Facebook AI, introduced a framework for learning latent features from raw audio using convolutional neural networks (CNNs) and self-supervised contrastive loss, achieving meaningful representations generalizable across tasks.
Wav2Vec 2.0 Baevski et al. (2020) improved on this with a masked prediction objective inspired by BERT Devlin et al. (2019), enabling better modelling of long-range dependencies in audio. With proper fine-tuning, its transformer-based architecture and contrastive loss enable state-of-the-art results in speech-processing tasks, especially emotion recognition Baevski et al. (2020).
Semantics.
Language models prioritize semantic understanding through contextual mechanisms. For instance, BERT employs bidirectional attention to consider both preceding and succeeding tokens, learning fine-grained contextual representations through a masked language modelling objective Devlin et al. (2019). Sentence-BERT (SBERT) extends BERT by mapping input sentences into dense vector spaces with a Siamese network, enabling semantic similarity comparisons via metrics like cosine similarity Reimers and Gurevych (2019).
However, SBERT is limited to textual data and cannot capture affective or prosodic nuances inherent to spoken language Mohebbi et al. (2021). This limitation underscores the need for multimodal approaches that integrate speech and text modalities to bridge the semantic and affective gaps in communication Zhang et al. (2023); Lian et al. (2023).
Multimodal Learning.
@@
Cross-modal alignment embeds data from different modalities into shared spaces to capture their relationships. Techniques like contrastive learning align related inputs (e.g., audio and text segments) while separating unrelated pairs Ye et al. (2022). Efforts to align text and audio include SpeechBERT Chuang et al. (2020), which adapted BERT’s framework to paired speech-text data, and SLAM (Speech-Language Aligned Models) Bapna et al. (2022), which optimized joint embedding spaces to improve downstream tasks like speech recognition and audio-text retrieval. Similarly, models like HuBERT Hsu et al. (2021) have shown promise in bridging text and audio through hierarchical feature learning. However, many models struggle with capturing prosodic and affective nuances, often focusing heavily on semantic alignment.
Figure 2: Diagram of WhiSA and WhiSPA training procedure involving a student-teacher model paradigm. Whisper (left) is semantically aligned to the ground truth embeddings encoded by SBERT (right). When PsychLex’s dimensions are included in the alignment function, the WhiSPA framework semantically and psychologically aligns the corresponding embeddings with contrastive loss criteria.
@@ 3 Data & Tasks
We utilize two psychological, mental health-focused datasets for training and evaluation: WTC-Segments (WTC) Kjell et al. (2024) and HiTOP-Segments (HiTOP) Kotov et al. (2022). WTC recordings were completed by patients in a clinic for WTC responders who came for a health monitoring visit. HiTOP interviews were completed by outpatients with psychiatric diagnoses who were recruited by the study team to complete a research interview. Both datasets consist of paired audio-text data, ensuring alignment between spoken content and its corresponding textual transcription.
@@
Total Segment Duration (min) 15,087 28,460
@@
From its source, WTC was curated from $\sim$6 minute interview recordings, on average, of patients responding to both personal and general questions in a structured manner Kjell et al. (2024). Contrarily, HiTOP followed a semi-structured format, where patients described experiences on set topics while also organically conversing with the interviewer. Once filtered for audio segments solely spoken by patients, interviews generally ranged from 45 to 90 minutes, yielding a voluminous and broadened set of audio segments Kotov et al. (2022). The recordings were diarized using NVIDIA NeMo and transcribed with openai/whisper-large-v2.
@@
To evaluate the encoding ability of WhiSPA for any given audio segment, we manually annotate a small subset from both datasets for valence and arousal dimensions expressed in their speech. Three random audio segments containing more than 5 uttered words from each user were sampled from each dataset and were annotated by two individuals with a background in psychology using the affective circumplex scale. This resulted in 300 audio segments, equally split between the two datasets.
@@ 4 Methodology
We begin with whisper-tiny pre-trained encoder-decoder model Radford et al. (2022) as our initial checkpoint prior to alignment. As seen in the Whisper (Student) portion of Figure 2, once the last non-padding token is predicted (<EOT>), we apply a mean pooling layer to the last hidden state of Whisper’s decoder yielding a singular representation for the input audio. This representation is then pooled using a learnable dense layer, and the output serves as our embedding during alignment. This aggregated representation is aligned to the pooled representations from pre-trained SBERT for semantic alignment and the PsychEmb’s dimensions for psychological alignment. We denote the pre-trained Whisper results in the paper with Whisper-384, referring to the number of hidden dimensions.
@@ 4.1 Alignment Objective
The alignment objective aims to improve the semantic and psychological information encoded in Whisper (student) with the help of the representations from a strong text encoding teacher model like SBERT and PsychEmb. In this work, we explore two suitable candidate objective functions to align speech representations with text, which are described below in detail.
@@ 4.1.1 Cosine Similarity Loss (CS)
The success of the cosine similarity-based approach in building geometrically robust representations in SBERT motivated its use as an alignment objective in this work. We apply cosine similarity loss to the pooled audio embeddings and pooled SBERT embeddings [^2] SBERT version: all-MiniLM-L12-v2, given by the following equation:
@@ 4.1.2 Noise Contrastive Estimation Loss (NCE)
The Noise Contrastive Loss (Equation 2) is optimized to increase the cosine similarity between a pair of audio embedding and its corresponding text embedding while simultaneously increasing the differentiation between the audio embedding and randomly sampled text embedding in that batch Ye et al. (2022).
@@
where $\mathcal{L}_{i}^{NCE}$ refers to contrastive loss criteria in which pairwise cosine similarities are calculated for each audio embedding with all text embeddings in that batch. Hence, there is only one positive text embedding that pairs with an audio embedding, while the remaining text embeddings from the batch serve as contrastive samples. Let $B(i)\in\mathcal{I}$, where $B(i)$ represents all other SBERT text embeddings in the batch such that $\mathbf{T}_{b}\neq\mathbf{T}_{i}$ Ye et al. (2022); Chen et al. (2020); Khosla et al. (2021). The variable $\mathbf{T}_{b}$ denotes the index of an arbitrary, negative SBERT text embedding sample and $\tau$, temperature, represents a trainable scalar parameter which is set to a default of $0.1$.
@@ 4.3 Adding Psychological Alignment (WhiSPA)
WhiSPA extends the WhiSA framework by augmenting PsychEmb dimensions into Whisper’s. While maintaining the semantic alignment objective, WhiSPA injects the PsychEmb dimensions into the SBERT embeddings under two settings: (1) with replacement: We adopted a naive strategy of replacing the first ten dimensions of SBERT’s embedding with the PsychEmb dimensions to maintain the same number of latent dimensions between both models. We use WhiSPA-384r to refer to this. (2) with projection: We concatenate the PsychEmb dimensions to the text embedding from SBERT. Consequently, this requires a $384\times 10$ learnable projection matrix, $P$, to transform Whisper embeddings of dimensionality $384$ to $394$, which is then transformed using a $TanH$ activation. This model goes by the name WhiSPA-394. To address the numerical instability issues from modelling the PsychEmb dimensions in its absolute range, we standardize and scale them to match their distribution to that of SBERT’s embeddings. Refer to Appendix subsection A.2 for more information on training.
Dataset Speech Model Traits States Dispositions OPE CON EXT AGR NEU VAL ARO ANG ANX DEP $r(\uparrow)$ $mse(\downarrow)$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ HiTOP Wav2Vec-768 .61 .15 .60 .15 .59 .13 .47 .11 .59 .14 .40 .001 .49 .001 .34 .04 .42 .02 .48 .04 Whisper-384 .74 .11 .80 .08 .69 .10 .76 .06 .78 .08 .71 .001 .82 .000 .53 .03 .61 .01 .65 .03 WhiSA-384 .71* .11 .81* .08 .70* .10 .77* .06 .78* .08 .73* .001 .83* .000 .59$\dagger$ .03 .61 .01 .61$\dagger$ .04 WhiSPA-384r .74* .11 .83$\dagger$ .07 .70* .10 .79$\dagger$ .05 .79$\dagger$ .07 .78$\dagger$ .000 .85$\dagger$ .000 .59$\dagger$ .03 .61$\dagger$ .01 .66* .03 WhiSPA-394 .72* .11 .83$\dagger$ .07 .72* .09 .79$\dagger$ .05 .82$\dagger$ .07 .76$\dagger$ .000 .84* .000 .62$\dagger$ .03 .65$\dagger$ .01 .63* .03 WTC Wav2Vec-768 .26 .56 .45 .59 .21 .69 .31 .48 .25 .68 .24 .004 .47 .006 .14 .23 .00 .16 .17 .16 Whisper-384 .57 .43 .70 .37 .68 .38 .64 32 .67 .40 .56 .003 .82 .002 .54 .16 .46 .13 .45 .13 WhiSA-384 .70$\dagger$ .31 .82$\dagger$ .24 .75$\dagger$ .32 .76$\dagger$ .23 .77$\dagger$ .30 .67$\dagger$ .002 .85$\dagger$ .002 .66$\dagger$ .13 .61$\dagger$ .10 .61$\dagger$ .10 WhiSPA-384r .71$\dagger$ .29 .82$\dagger$ .24 .74$\dagger$ .30 .76$\dagger$ .20 .76$\dagger$ .27 .68$\dagger$ .002 .85$\dagger$ .002 .67$\dagger$ .01 .61$\dagger$ .09 .61$\dagger$ .09 WhiSPA-394 .72$\dagger$ .28 .83$\dagger$ .22 .76$\dagger$ .29 .79$\dagger$ .19 .79$\dagger$ .26 .70$\dagger$ .002 .86$\dagger$ .002 .69$\dagger$ .11 .64$\dagger$ .09 .66$\dagger$ .09

### v2

@@ Abstract
Current speech encoding pipelines often rely on an additional text-based LM to get robust representations of human communication, even though SotA speech-to-text models often have a LM within. This work proposes an approach to improve the LM within an audio model such that the subsequent text-LM is unnecessary. We introduce WhiSPA (Whisper with Semantic and Psychological Alignment), which leverages a novel audio training objective: contrastive loss with a language model embedding as a teacher. Using over 500k speech segments from mental health audio interviews, we evaluate the utility of aligning Whisper’s latent space with semantic representations from a text autoencoder (SBERT) and lexically derived embeddings of basic psychological dimensions: emotion and personality. Over self-supervised affective tasks and downstream psychological tasks, WhiSPA surpasses current speech encoders, achieving an average error reduction of $73.4\%$ and $83.8\%$, respectively. WhiSPA demonstrates that it is not always necessary to run a subsequent text LM on speech-to-text output in order to get a rich psychological representation of human communication.
@@
Rajath Rao1, Adithya V Ganesan1, Oscar Kjell1, Jonah Luby1, Akshay Raghavan1, Scott Feltman1 Whitney Ringwald2, Ryan L. Boyd3, Benjamin Luft1, Camilo Ruggero3, Neville Ryant4, Roman Kotov1, H. Andrew Schwartz1 1Stony Brook University, 2University of Minnesota 3University of Texas at Dallas, 4University of Pennsylvania Correspondence: rajath.rao@stonybrook.edu, has@cs.stonybrook.edu
@@ 1 Introduction
Figure 1: Speech processing pipelines that use text-based representations from language models often yield higher accuracies than those produced solely by SotA audio models. While both approaches ingest the same raw audio signal, we close the performance gap by introducing a speech encoder with similar performance to a text-based LM pipeline.
Human communication is inherently multimodal, but AI integration of modalities is often fragmented Lazaro et al. (2021); Gu et al. (2017), where speech models, such as Whisper Radford et al. (2022), are often pipelined into text-based language models (LMs) Chuang et al. (2020) in order to get the most accurate speech-based representations (see Figure 1). Text-based LMs produce richer semantic representations Wu et al. (2024); Fu (2024). This often results in redundant computational costs from having two LMs in the pipeline (one within the audio model and one for the text LM) and representations remain incomplete of the full spectrum of human expressions Zhang et al. (2023); Lian et al. (2023). y important for psychological and social scientific applications of LMs where reptions from text-based LMs demonstrate superior performance than direct speech repdirect speechchen2024textbetteLukac2024-tb; rspeech.
@@
Here, we seek to bridge the semantic and psychological representation gap between speech-based LMs present in audio models and text-based LMs. We introduce a speech encoding model, WhiSPA [^1] https://github.com/humanlab/WhiSPA (Whisper with Semantic and Psychological Alignment), which aligns a pre-trained speech recognition model, Whisper Radford et al. (2022), with the latent dimensions from SBERT Reimers and Gurevych (2019), intended to better capture semantics and deeper psychological information V Ganesan et al. (2022); Park et al. (2014). Such alignment reduces computational and memory inefficiencies, circumventing the need for a second text encoder, as it enables a unified cross-modal representation between speech and language models. Still, since text is derivable from speech, speech should intrinsically be mappable to the same rich semantic features from the text.
Our focus on psychological or human-level tasks reflects a growing demand for foundation models to better understand the intrinsic qualities of human communication Soni et al. (2024). As  Clark and Schober (1992) put it, “The common misconception is that language has to do with words and what they mean. It does not. It has to do with people and what they mean.” and specifically how well the representations capture information about the people communicating Hovy and Yang (2021); Soni et al. (2022). More specifically, psychological studies have suggested mental health attributes are highly multimodal as they are influenced by subtle nuances in voice and content Sartori and Orrù ; Chen et al. (2024).
Our main contributions include: (1) The development of WhiSPA (Whisper with Semantic and Psychological Alignment), with a novel alignment objective, (2) Evaluation of the hypothesis that aligning text and audio latent spaces can significantly enhance audio-based representations for a deeper semantic and psychological understanding of human communication, (3) Demonstration of significant accuracy improvements in self-supervised tasks and downstream psychological tasks over systematically tested variants of WhiSPA. We find that: (a) aligning with text-based semantic and psychological representations drastically improves audio representations, including SotA person-level psychological assessments; (b) a Noise Contrastive Estimation loss yielded a more optimal convergence in aligning Whisper’s latent space with semantic and psychological dimensions. and (c) for downstream psychological tasks, there was almost no benefit in utilizing SBERT representations on top of WhiSPA’s, suggesting the same information from a text LM can be captured with the LM of the audio model and thus it is not necessary to pipeline another text LM after the audio model.
@@ 2 Background
This work builds on top of Whisper Radford et al. (2022), OpenAI’s SotA automatic speech recognition (ASR) foundation model. We chose Whisper over other alternatives such as HuBERT and Wav2Vec2-BERT, since previous works Kyung et al. (2024); Yang et al. (2023) have shown that Whisper has a stronger language encoding module at capturing speaker attributes.
Recent advances in foundational speech technologies, like Whisper and HuBERT, have vastly improved the performances on speech recognition tasks Radford et al. (2022); Hsu et al. (2021). However, they have limited ability to capture deeper semantics and speaker attributes compared to a text-based language model Chen et al. (2024); Dong et al. (2022). Prior works that have addressed this have targeted a very narrow scope of psychological attributes Busso et al. (2008). These gaps underscore the need for methodologies that bridge speech encoders’ acoustic robustness with the psychological depth of text-based language models—a challenge we address by embedding fundamental psychological dimensions present in one’s speech.
Multi-level fusion architectures leveraging both acoustic and lexical features have shown to improve the performance on downstream tasks. For instance, Zhao et al. (2022) demonstrates that coattention-based early fusion and late fusion using Wav2Vec2.0 Baevski et al. (2020); Schneider et al. (2019) and BERT Devlin et al. (2019) outperform SotA emotion recognition benchmarks.Other recent works inject acoustic nuances into language models using textual descriptions of speech characteristics Wu et al. (2024) or common-sense reasoning through historical utterances from the speaker Fu (2024).However, this approach does not fully leverage the cross-modal dependencies between text and audio, as it remains unimodal, relying solely on textual inputs rather than raw acoustic representations.
@@
Prior works in cross-modal alignment provide foundational insights for this integration. Compositional Contrastive Learning Chen et al. (2021) distilled audio-visual knowledge into video representations by aligning teacher-student embeddings across modalities, embedding rich semantics from teacher-audio and image models into the student-video model. In another work,  Dong et al. (2022) improved the accuracy of intent classification of spoken language by employing a contrastive loss using both speech and language features. These works highlight that the cross-modal alignment objective embeds information from different modalities into shared spaces to capture their relationships, while contrastive learning aids in grouping related inputs across different modalities (e.g., audio and text segments) while separating unrelated pairs Ye et al. (2022). Efforts to align text and audio include SpeechBERT Chuang et al. (2020), which adapted BERT’s framework Devlin et al. (2019) to paired speech-text data, and SLAM (Speech-Language Aligned Models) Bapna et al. (2022), which optimized joint embedding spaces to improve downstream tasks like speech recognition and audio-text retrieval. To the best of our knowledge, this is the first work to perform cross-modal learning to endow the foundational speech model with richer semantic and psychological representations.
@@ 3 Data & Tasks
We utilize two psychological, mental health-focused datasets for training and evaluation: WTC-Segments (WTC) Kjell et al. (2024) and HiTOP-Segments (HiTOP) Kotov et al. (2022). WTC recordings were completed by patients in a clinic for World Trade Center (9/11) responders who came for a health monitoring visit. HiTOP interviews were completed by outpatients with psychiatric diagnoses who were recruited by the study team to complete a research interview. Both datasets consist of paired audio-text data, ensuring alignment between spoken content and its corresponding textual transcription.
Figure 2: Diagram of WhiSA and WhiSPA training procedure involving a student-teacher model paradigm. Whisper (left) is semantically aligned to the ground truth embeddings encoded by SBERT (right). When PsychEmb features are included in the alignment function, the WhiSPA framework semantically and psychologically aligns the corresponding embeddings with contrastive loss criteria.
@@
Total Segment Duration (hr) $\sim$252 $\sim$474
@@
From its source, WTC was curated from $\sim$6 minute interview recordings, on average, of patients responding to both personal and general questions in a structured manner Kjell et al. (2024). Contrarily, HiTOP followed a semi-structured format, where patients described experiences on set topics while also organically conversing with the interviewer. Once filtered for audio segments solely spoken by patients, interviews generally ranged from 45 to 90 minutes, yielding a voluminous and broadened set of audio segments Kotov et al. (2022). The recordings were diarized using NVIDIA NeMo and transcribed with whisper-large-v2.
@@
To evaluate the encoding ability of WhiSPA for any given audio segment, we manually annotate a small subset from both datasets for valence and arousal dimensions expressed in their speech. Three random audio segments containing more than 5 uttered words from each user were sampled from each dataset and were annotated by two individuals with a background in psychology using the affective circumplex scale (Figure 7). This resulted in 300 audio segments, equally split between the two datasets.
@@ 4 Methodology
Aligning audio representations directly with a text-based language model allows us to infuse the audio model’s latent space with the rich semantic and affective details typically provided by text representations, thereby eliminating the need for a separate text LM. While this approach does not explicitly leverage the unique acoustic features of speech, it prioritizes efficiency by avoiding redundant processing and consistently delivers a semantically enriched representation—an advantage that is particularly critical for psychological and social scientific applications Lukac (2024); Chen et al. (2024).
@@
We begin with the Whisper [^2] Whisper-384 version: whisper-tiny encoder-decoder backbone Radford et al. (2022), which does not run autoregressively. During training, audio segments are previously transcribed with whisper-large-v2, making it entirely self-supervised. Likewise, SBERT and PsychEmb representations were encoded using these transcriptions. As seen in the Whisper (Student) portion of Figure 2, we apply a mean pooling layer to the last hidden state of Whisper’s decoder yielding a singular representation for the input audio. This representation is then pooled using a learnable dense layer, and the output serves as our embedding during alignment. This aggregated representation is aligned to the pooled representations from pre-trained SBERT for semantic alignment and the PsychEmb’s dimensions for psychological alignment. Throughout this paper, we denote the pre-trained Whisper model as Whisper-384, where the numeric suffix refers to the embedding dimensionality.
@@ 4.1 Alignment Objective
While fusion architectures focus on merging acoustic-textual features throughout layers, we contrast this paradigm by directly aligning cross-modal latent spaces for deeper semantic and psychological representations from audio, bypassing the need for task-specific fusion architectures. Our alignment objective aims to improve the semantic and psychological information encoded in Whisper (student) with the help of the representations from a strong text encoding teacher model like SBERT [^3] SBERT-384 version: all-MiniLM-L12-v2 and PsychEmb. In this work, we explore two suitable candidate objective functions to align speech representations with text, which are described below in detail.
@@ 4.1.1 Cosine Similarity Loss (CS)
The success of the cosine similarity-based approach in building geometrically robust representations in SBERT motivated its use as an alignment objective in this work. We apply cosine similarity loss to the pooled audio embeddings and pooled SBERT embeddings, given by the following equation:
@@ 4.1.2 Noise Contrastive Estimation Loss (NCE)
The Noise Contrastive Loss (Equation 2) is optimized to increase the cosine similarity between a pair of audio embedding and its corresponding text embedding while simultaneously increasing the differentiation between the audio embedding and randomly sampled text embeddings in that batch Ye et al. (2022).
@@
where $\mathcal{L}_{i}^{NCE}$ refers to contrastive loss criteria in which pairwise cosine similarities are calculated for each audio embedding with all text embeddings in that batch. Hence, there is only one positive text embedding that pairs with an audio embedding, while the remaining text embeddings from the batch serve as contrastive samples. Let $B(i)\in\mathcal{I}$, where $B(i)$ represents all other SBERT text embeddings in the batch such that $\mathbf{T}_{b}\neq\mathbf{T}_{i}$ Ye et al. (2022); Chen et al. (2020); Khosla et al. (2021). The variable $\mathbf{T}_{b}$ denotes the index of an arbitrary, negative SBERT text embedding sample and $\tau$, temperature, represents a tunable scalar parameter which is set to $0.1$.
@@ 4.3 Adding Psychological Alignment (WhiSPA)
WhiSPA extends the WhiSA framework by augmenting PsychEmb dimensions into Whisper’s. While maintaining the semantic alignment objective, WhiSPA injects the PsychEmb dimensions into the SBERT embeddings under two settings: (1) with replacement: We adopted a naive strategy of replacing the first ten dimensions of SBERT’s embedding with the PsychEmb dimensions to maintain the same number of latent dimensions between both models. We use WhiSPA-384r to refer to this. (2) with projection: We concatenate the PsychEmb dimensions to the text embedding from SBERT. Consequently, this requires a $384\times 10$ learnable projection matrix, $P$, to transform Whisper embeddings of dimensionality $384$ to $394$, which is then passed through a $TanH$ activation. This model goes by the name WhiSPA-394. To address the numerical instability issues from modeling the PsychEmb dimensions in its absolute range, we standardize and scale them to match SBERT’s distribution of embedding values. Refer to Appendix subsection A.2 for more information on training.
Dataset Model Traits States Dispositions OPE CON EXT AGR NEU VAL ARO ANG ANX DEP $r(\uparrow)$ $mse(\downarrow)$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ HiTOP W2V2B .63 .14 .69 .12 .75 .09 .60 .09 .72 .10 .65 .001 .73 .000 .45 .04 .51 .02 .64 .03 HuBERT .67 .13 .71 .11 .77 .08 .57 .10 .70 .11 .66 .001 .73 .000 .48 .04 .48 .02 .58 .04 Whisper-384 .74 .11 .80 .08 .69 .10 .76 .06 .78 .08 .71 .001 .82 .000 .53 .03 .61 .01 .65 .03 SBERT-384 .73 .11 .83 .07 .68 .11 .75 .06 .77 .09 .69 .001 .81 .000 .59 .03 .60 .01 .61 .04 WhiSA-384 .71* .11 .81* .08 .70 .10 .77* .06 .78* .08 .73* .001 .83* .000 .59$\dagger$ .03 .61 .01 .61 .04 WhiSPA-384r .74* .11 .83$\dagger$ .07 .70 .10 .79$\dagger$ .05 .79$\dagger$ .07 .78$\dagger$ .000 .85$\dagger$ .000 .59$\dagger$ .03 .61$\dagger$ .01 .66* .03 WhiSPA-394 .72* .11 .83$\dagger$ .07 .72 .09 .79$\dagger$ .05 .82$\dagger$ .07 .76$\dagger$ .000 .84* .000 .62$\dagger$ .03 .65$\dagger$ .01 .63* .03 WTC W2V2B .33 .54 .51 .55 .34 .64 .37 .46 .34 .64 .32 .004 .51 .005 .31 .21 .14 .16 .22 .15 HuBERT .35 .54 .57 .50 .39 .61 .44 .43 .42 .60 .38 .003 .53 .005 .36 .20 .15 .16 .22 .16 Whisper-384 .57 .43 .70 .37 .68 .38 .64 32 .67 .40 .56 .003 .82 .002 .54 .16 .46 .13 .45 .13 SBERT-384 .65 .35 .78 .29 .73 .33 .73 .25 .73 .34 .62 .003 .86 .002 .62 .14 .56 .11 .59 .11 WhiSA-384 .70$\dagger$ .31 .82$\dagger$ .24 .75$\dagger$ .32 .76$\dagger$ .23 .77$\dagger$ .30 .67$\dagger$ .002 .85$\dagger$ .002 .66$\dagger$ .13 .61$\dagger$ .10 .61$\dagger$ .10 WhiSPA-384r .71$\dagger$ .29 .82$\dagger$ .24 .74$\dagger$ .30 .76$\dagger$ .20 .76$\dagger$ .27 .68$\dagger$ .002 .85$\dagger$ .002 .67$\dagger$ .01 .61$\dagger$ .09 .61$\dagger$ .09 WhiSPA-394 .72$\dagger$ .28 .83$\dagger$ .22 .76$\dagger$ .29 .79$\dagger$ .19 .79$\dagger$ .26 .70$\dagger$ .002 .86$\dagger$ .002 .69$\dagger$ .11 .64$\dagger$ .09 .66$\dagger$ .09